OpenStudy (zzr0ck3r):

Prove that in Hilbert spaces \((A^{\perp})^{\perp}=\overline{\text{span}(A)}\).

OpenStudy (zzr0ck3r):

Im not quite sure what I need to show here @eliassaab

OpenStudy (zzr0ck3r):

So if we are in \(\mathbb{R}^2\) and we take some line through the origin, and we take some vector that is orthogonal to that line and call that vector b and set A={b}, then the orthogonal complement of A is all the vectors that make up the entire line and the orthogonal complement of that is all the linear combinations of b, or in this case, ab where a is in R. Am I thinking of this right?

OpenStudy (anonymous):

It is clear \[ x\in(B^{\perp})^{\perp}\cap B^{\perp}=\{ 0 \}\\ <x,x>=0=||x||^2\implies x=0 \]

OpenStudy (anonymous):

Let G be a closed subspace of H then\[ (G^{\perp})^{\perp}= G \] All we need to prove that \[ (G^{\perp})^{\perp}\subset G \] Now \[ H= G \oplus G^{\perp} \\\text{ Let } x\in (G^{\perp})^{\perp}\ \\ x= s+ t; \quad s\in G, t\in G^{\perp}\\ t= x-s \in (G^{\perp})^{\perp} \cap G^{\perp}=\{0\}\\ x=s \in G \] We are done for that part

OpenStudy (anonymous):

Let finish your problem Let\[ G= \overline{\text{span}(A)} \] It is easy to see that \( G \subset (A^{\perp})^{\perp}\) \[ A \subset G \\ G^{\perp} \subset A^{\perp}\\ (A^{\perp})^{\perp}\subset (G^{\perp})^{\perp}=G \] We are done.

OpenStudy (anonymous):

Let G be a closed subspace of a Hilbert space H, let h be an element in H not in G, then there exist a unique element t in G so that d(h,G)=|| h-t|| By definition one can write that \[ d(h,G)\le ||h -t_n|| \le d(h,G) + \frac 1 n \] It is clear that |( t_n |) is Cauchy and its limit t satisfies d(h,G)=|| h-t|| Uniqueness. Suppose that there are s and t so that k=d(h,G)=|| h-t||=||h-s|| Write \[ 0\le || s-t||^2 =|| s- h + h-t||^2= 2||s-h||^2 + 2||h-t||^2 - ||s-h -(h-t)||^2=\\ 2||s-h||^2 + 2||h-t||^2 - ||s+t -2 h||^2=\\ 2||s-h||^2 + 2||h-t||^2 - ||2 \frac{s+t}2 -2 h||^2=\\ 2||s-h||^2 + 2||h-t||^2 - 4|| \frac{s+t}2 - h||^2\le\\ 2 k^2+2 k^2 -4 k^2=0\\Hence\\ s=t \] This element h-t is perpendiculat to every element in G. To see why, write \[ z=h-t\\\ g \in G\ \\ ||z||^2 = || h-t||^2 \le || h - ( t+ \frac{<z,g> g}{||g||^2} ||^2=\\ || h - t - \frac{<z,g> g}{||g||^2} ||^2=|| z - \frac{<z,g> g}{||g||^2} ||^2=\\ ||z||^2 -2 \frac{|<z,g>|^2}{||g||^2} +\frac{|<z,g>|^2}{||g||^2}=\\ ||z||^2 - \frac{|<z,g>|^2}{||g||^2} \implies <z,g>=0 \]

OpenStudy (anonymous):

It is easy to show now that if G is closed subspace of H then \[ H=G \oplus G^{\perp} \] Try it

OpenStudy (zzr0ck3r):

great, ill go play with this and get back to you if I have any questions. thanks again

OpenStudy (anonymous):

YW